Crane Company, organized in 2019, has set up a single account for all intangible assets. The following summary discloses theMake the entry as of December 31, 2020, recording any necessary amortization. (Credit account titles are automatically indentReflect all balances accurately as of December 31, 2020. Use straight-line amortization ta ta ta ta
Answer #1
1
Account Titles and Explanation Debit Credit
Patents 407700
Goodwill 363000
Franchises 443000
Copyrights 171600
Research and Development Expense 233000
         Intangible Assets 1618300
2
Account Titles and Explanation Debit Credit
Amortization Expense 81750
      Patents 45300 =407700/9
      Franchise 22150 =443000/10*6/12
      Copyright 14300 =171600/5*5/12
3
Use straight-line amortization
Patents 362400 =407700-45300
Goodwill 363000
Franchise 420850 =443000-22150
Copyright 157300 =171600-14300